1.0 mol of gas at a temperature of -120 ∘C fills a 3.3 L container. Part A What is the gas pressure?
Answer #1

here,

number of moles , n = 1

temprature , T = - 120 degree C

T = 153 K

volume , V = 3.3 L

R = 0.0821

let the pressure be P

using ideal gas equation

P*V = n * R * T

P * 3.3 = 1 * 0.0821 * 153

P = 3.81 atm

the gas pressure is 3.81 atm